# ORION: Option-Regularized Deep Reinforcement Learning for Cooperative Multi-Agent Online Navigation

### v1

@@ I Introduction
Early works in multi-agent navigation primarily focused on ensuring collision-free motion at the kinematic level, typically by decoupling target-driven planning from local obstacle/neighboring robots avoidance [jin2021hierarchical, yu2023learning]. These approaches largely treat interactions among the robot team as safety constraints rather than opportunities to enhance team-level performance: robots rarely leverage team-level coordination and shared global information to guide long-horizon decision-making. Another line of work studies formation-preserving or centroid-driven navigation, where coordinated motions yield more globally consistent behaviors, e.g., maintaining formation structure and avoiding oscillations/conflicts [tanner2012multiagent, quan2023robust]. Yet, these approaches typically assume a known environment or fixed/specific formation objectives, which limits their flexibility in open-ended navigation tasks where robots must dynamically adapt their behaviors to evolving map information. Another closely related domain is Multi-Agent Path Finding (MAPF), where the goal is to compute globally coordinated, conflict-free trajectories for all robots (usually offline), often emphasizing scalability in large structured environments [li2021eecbs, li2022mapf, sartoretti2019primal]. Despite the global coordination achieved by MAPF approaches, they usually assume a fully known environment and lack the ability to reason about partial observability, exploration, or online map updates.
@@ II-A Navigation in Partially Known Environments
Navigation in fully known environments has been extensively studied, with search-based planners relying on efficient heuristic-guided graph expansion and sampling-based planners exploring the continuous space through incremental sampling [koenig2002d, likhachev2005anytime, karaman2011sampling]. While these methods are effective in static and fully known settings, they tend to degrade in partially known environments due to their limited ability to anticipate unobserved environment and the heavy computation required for frequent replanning.
Learning-based navigation methods, particularly deep reinforcement learning, have shown promise in reasoning over uncertainty without hand-crafted heuristics. Many of these approaches, however, are designed primarily for local/motion-level planning with sensory inputs such as RGB-D images or point clouds [pfeiffer2017perception, jin2020mapless], making them difficult to scale to large, complex environments where long-horizon reasoning is essential [tai2017virtual]. Several approaches employ belief-informed [liang2023context] or hierarchical policies [liang2024hdplanner] to enhance spatial reasoning and long-term waypoint planning, but still require additional mechanisms to prune redundant exploration or maintain plan quality at scale. Parallel efforts incorporate generative models to predict unknown regions [ho2025mapex, wang2025cogniplan] or to guide long-term planning [cao2025dare], which alleviates myopic behaviors. Overall, although these methods significantly advance navigation under uncertainty, they operate as heuristic, single-agent policies and do not benefit from team-level cooperation. Moreover, their decision-making is tightly coupled to individual belief updates and is not suitable to handle inter-agent interactions, making them difficult to transfer directly to multi-agent navigation settings.
@@ II-B Cooperative Multi-Agent Navigation
Beyond single-robot navigation, achieving cooperation across a team of robots poses more challenges. Most works in multi-agent navigation focus on guaranteeing collision-free motion at the kinematic level [jin2021hierarchical, yu2023learning], often in cluttered/crowded environments. These methods primarily treat other agents as local dynamic constraints, and their policies are designed to ensure short-horizon safety and feasibility. While some of them explicitly consider environmental uncertainty [wagner2017path], this decision-making is still tightly coupled to individual belief updates, making it difficult to endow robots with long-horizon, team-level reasoning in large, partially known environments. Another line of work studies multi-agent navigation while keeping a free-floating formation of specific shape [tanner2012multiagent, quan2023robust]. These approaches design behavior-based rules to induce globally consistent motion, but they typically assume known environments, fixed formation objectives, or pre-specified connectivity patterns.
Multi-Agent Path Finding (MAPF) offers a more global view of coordination, aiming to compute conflict-free trajectories that minimize makespan. Representative methods include bounded-suboptimal [li2021eecbs], prioritized/rule-based search [li2022mapf], as well as learning-augmented [sartoretti2019primal] and lifelong variants [jiang2025deploying] that scale to thousands of agents. Despite their strong scalability and coordination in discrete, fully known environments, extending MAPF to continuous, partially known environments with online execution remains challenging. A few works move closer to cooperative reasoning by studying multi-agent path topologies without explicit communication [mavrogiannis2019multi], enabling targeted rendezvous/message passing among agents [das2019tarmac, tan2024ir], or co-optimizing policies with reconfigurable environments [gao2025co]. Yet, they remain largely heuristics-based, relying on implicit communication that emerges from local observations rather than deliberate intent sharing, and thus lack mechanisms to model temporally extended, altruistic cooperation, such as when to sacrifice individual progress or re-engage to assist teammates.
@@ III-B Observation as Informative Graph
Each node/viewpoint is augmented with semantic attributes beyond 2D coordinates to obtain richer representations: (1) Prior utility $u_{p,j}$: extending the notion of frontiers [yamauchi1998frontier], we define prior frontiers $\mathcal{F}_{p}$ as boundaries between free space and uncertain obstacles, and let $u_{p,j}$ denote the number of such prior frontiers that are visible, i.e., along a collision-free line of sight, from $\hat{v}_{j}\in\hat{V}_{t}$ within a sensing range $r_{\rm fov}$. (2) Utility $u_{j}$: defines the same quantity but computes it on the current map $\mathcal{M}$. (3) Visited flag $\delta_{i,j}$: indicates whether agent $i$ has visited $v_{j}$. (4) Verified signal $s_{j}$: marks whether $v_{j}$ lies within $\mathcal{M}$. (5) Occupancy indicator $p_{j}$: encodes, as a ternary variable, whether $v_{j}$ is currently occupied by agent $i$, by another agent, or by no agent. (6) Target binary $t_{j}$: specifies whether $v_{j}$ coincides with agent $i$’s own target or with another agent’s target (if any). We further cluster all viewpoints with non-zero utility within a range $r_{b}$ into a beacon set [liang2024hdplanner], representing candidate regions to navigate. We further construct a planning graph $\bar{G}_{t}=(\bar{V}_{t},\bar{E}_{t})$ on $\mathcal{M}$, containing samples from both free and unknown cells, where edges connect collision-free pairs of neighbors. These structures allow us to define: (7) Navigation guidepost $g_{n,j}$: the Dijkstra trajectory on $\bar{G}_{t}$ from $v_{j}$ to the beacon nearest the agent’s assigned target. (8) Cooperation guidepost $g_{c,j}$: if any target remains unverified, i.e., not in the explored area, we compute the trajectory from $v_{j}$ to the beacon nearest the closest unverified target; otherwise $g_{c,j}=g_{n,j}$ if no targets remain.
@@ III-D2 Option Decoder $\pi_{\vartheta}$
During multi-agent navigation, an agent may either navigate toward its designated target or assist others by collecting more information that benefits the team, each demanding distinct and sometimes conflicting reasoning. In such cases, adaptive switching between self-directed and cooperative behaviors becomes crucial. To balance individual efficiency and collective performance for improved overall task success, we introduce the option-critic structure [bacon2017option], where an option denotes a high-level behavioral mode characterized by a termination head function and an intra-option policy, while the critic provides a value-based estimation guiding high-level option/low-level action optimization. Specifically, we augment the policy network with an option decoder, which determines whether to continue the current high-level behavior or terminate and switch to another.
@@
MAContext w/o dual-stage [liang2023context] 522.07 ($\pm 162.09$) 64.59 510.11 ($\pm 170.3$) 62.64 495.39 ($\pm 151.08$) 61.03 583.66 ($\pm 177.65$) 71.52
MAContext [liang2023context] 503.3 ($\pm 152.98$) 61.63 495.45 ($\pm 129.99$) 61 478.49 ($\pm 110.82$) 59.16 500.77 ($\pm 168.78$) 60.53
EECBS [li2021eecbs] 526.12 ($\pm 151.22$) 66.08 501.78 ($\pm 127.77$) 62.64 477.12 ($\pm 129.86$) 59.49 468.3 ($\pm 136.55$) 58.6
LNS2 [li2022mapf] 520.99 ($\pm 141.15$) 65.09 499.24 ($\pm 119.46$) 62.52 474.32 ($\pm 127.82$) 59.49 466.24 ($\pm 141.68$) 58.44
@@ III-E1 Multi-Agent Centralized Critic
Following MAAC [iqbal2019actor], the critic computes a query $\mathbf{q}_{i}=W_{q}\mathbf{h}_{i}$ and attends to keys and values $\mathbf{k}_{j}=W_{k}\mathbf{u}_{j}$, $\mathbf{v}_{j}=W_{v}\mathbf{u}_{j}$ via
@@ III-E2 Option-Regularized Critic Learning
Our critic network adopts the same graph-transformer backbone as the policy network, but differs in its use of a privileged graph encoder. During training, this encoder has access to privileged ground-truth information (i.e., the ground-truth map), enabling the critic to provide low-variance estimations on long-term returns to the decentralized policies. Unlike the actor’s dual encoders (one for the combined graph and one for the current graph), the critic employs a single unified encoder that captures the complete relational context of all agents and their robot beliefs. Besides the state representation input, the critic also takes the option feature selected at the previous step as input, which benefits both option and waypoint optimization during critic learning [iqbal2019actor]. The value estimate is explicitly conditioned on the agent’s previous option $z_{t-1}$, representing the high-level mode active at the current step. Specifically, the option embedding is added to the current node feature:
@@
On the other hand, following [haarnoja2018soft], the policy loss is further improved by our option-regularized policy network:
@@
Finally, the overall loss is $\mathcal{L}_{\text{policy}}=\mathcal{L}_{\pi}+\lambda_{\vartheta}\,\mathcal{L}_{\vartheta},$ where $\lambda_{\vartheta}$ balances termination learning. Following [haarnoja2018soft], the critic is optimized via a temporal-difference (TD) regression objective.
@@ IV-A Comparisons in Simulated Maps
Inspired by recent navigation and exploration benchmarks [liang2023context, tan2024ir], we introduce a new benchmark comprising hundreds of large, maze-like, and challenging simulated maps, in which we evaluate ORION against representative baselines under partially known environments. For the learning-based baseline, we consider MAContext [liang2023context], a multi-agent extension of a DRL navigation planner designed for partially known environments. For classical MAPF solvers, we include 1) LNS2 [li2022mapf], a highly efficient and scalable planner that iteratively repairs collision-containing path sets by replanning for a subset of colliding agents, and 2) EECBS [li2021eecbs], a bounded-suboptimal MAPF planner that combines Explicit Estimation Search at the high level with focal search at the low level. Since both LNS2 and EECBS operate on static maps, we deploy them in our partially known setting by triggering replanning whenever new changes are observed by any agent in the environment. To further analyze the contribution of our design, we additionally evaluate two ablation variants: 1) ORION w/o option, removing our option-critic structure, and 2) ORION w/o dual-stage, disabling our pre-/post-arrival stages. These variants help isolate the role of our proposed option-critic’s hierarchical decision-making and dual-stage cooperation strategy.

### v2

@@ I Introduction
Early works in multi-agent navigation primarily focused on ensuring collision-free motion at the kinematic level, typically by decoupling target-driven planning from local obstacle/neighboring robots avoidance [10, 28]. These approaches largely treat interactions among the robot team as safety constraints rather than opportunities to enhance team-level performance: robots rarely leverage team-level coordination and shared global information to guide long-horizon decision-making. Another line of work studies formation-preserving or centroid-driven navigation, where coordinated motions yield more globally consistent behaviors, e.g., maintaining formation structure and avoiding oscillations/conflicts [24, 20]. Yet, these approaches typically assume a known environment or fixed/specific formation objectives, which limits their flexibility in open-ended navigation tasks where robots must dynamically adapt their behaviors to evolving map information. Another closely related domain is Multi-Agent Path Finding (MAPF), where the goal is to compute globally coordinated, conflict-free trajectories for all robots (usually offline), often emphasizing scalability in large structured environments [14, 13, 21]. Despite the global coordination achieved by MAPF approaches, they usually assume a fully known environment and lack the ability to reason about partial observability, exploration, or online map updates.
@@ II-A Navigation in Partially Known Environments
Navigation in fully known environments has been extensively studied, with search-based planners relying on efficient heuristic-guided graph expansion and sampling-based planners exploring the continuous space through incremental sampling [12, 17, 11]. While these methods are effective in static and fully known settings, they tend to degrade in partially known environments due to their limited ability to anticipate unobserved environment and the heavy computation required for frequent replanning.
Learning-based navigation methods, particularly deep reinforcement learning, have shown promise in reasoning over uncertainty without hand-crafted heuristics. Many of these approaches, however, are designed primarily for local/motion-level planning with sensory inputs such as RGB-D images or point clouds [19, 9], making them difficult to scale to large, complex environments where long-horizon reasoning is essential [22]. Several approaches employ belief-informed [16] or hierarchical policies [15] to enhance spatial reasoning and long-term waypoint planning, but still require additional mechanisms to prune redundant exploration or maintain plan quality at scale. Parallel efforts incorporate generative models to predict unknown regions [6, 26] or to guide long-term planning [2], which alleviates myopic behaviors. Overall, although these methods significantly advance navigation under uncertainty, they operate as heuristic, single-agent policies and do not benefit from team-level cooperation. Moreover, their decision-making is tightly coupled to individual belief updates and is not suitable to handle inter-agent interactions, making them difficult to transfer directly to multi-agent navigation settings.
@@ II-B Cooperative Multi-Agent Navigation
Beyond single-robot navigation, achieving cooperation across a team of robots poses more challenges. Most works in multi-agent navigation focus on guaranteeing collision-free motion at the kinematic level [10, 28], often in cluttered/crowded environments. These methods primarily treat other agents as local dynamic constraints, and their policies are designed to ensure short-horizon safety and feasibility. While some of them explicitly consider environmental uncertainty [25], this decision-making is still tightly coupled to individual belief updates, making it difficult to endow robots with long-horizon, team-level reasoning in large, partially known environments. Another line of work studies multi-agent navigation while keeping a free-floating formation of specific shape [24, 20]. These approaches design behavior-based rules to induce globally consistent motion, but they typically assume known environments, fixed formation objectives, or pre-specified connectivity patterns.
Multi-Agent Path Finding (MAPF) offers a more global view of coordination, aiming to compute conflict-free trajectories that minimize makespan. Representative methods include bounded-suboptimal [14], prioritized/rule-based search [13], as well as learning-augmented [21] and lifelong variants [8] that scale to thousands of agents. Despite their strong scalability and coordination in discrete, fully known environments, extending MAPF to continuous, partially known environments with online execution remains challenging. A few works move closer to cooperative reasoning by studying multi-agent path topologies without explicit communication [18], enabling targeted rendezvous/message passing among agents [3, 23], or co-optimizing policies with reconfigurable environments [4]. Yet, they remain largely heuristics-based, relying on implicit communication that emerges from local observations rather than deliberate intent sharing, and thus lack mechanisms to model temporally extended, altruistic cooperation, such as when to sacrifice individual progress or re-engage to assist teammates.
@@ III-B Observation as Informative Graph
Each node/viewpoint is augmented with semantic attributes beyond 2D coordinates to obtain richer representations: (1) Prior utility $u_{p,j}$: extending the notion of frontiers [27], we define prior frontiers $\mathcal{F}_{p}$ as boundaries between free space and uncertain obstacles, and let $u_{p,j}$ denote the number of such prior frontiers that are visible, i.e., along a collision-free line of sight, from $\hat{v}_{j}\in\hat{V}_{t}$ within a sensing range $r_{\rm fov}$. (2) Utility $u_{j}$: defines the same quantity but computes it on the current map $\mathcal{M}$. (3) Visited flag $\delta_{i,j}$: indicates whether agent $i$ has visited $v_{j}$. (4) Verified signal $s_{j}$: marks whether $v_{j}$ lies within $\mathcal{M}$. (5) Occupancy indicator $p_{j}$: encodes, as a ternary variable, whether $v_{j}$ is currently occupied by agent $i$, by another agent, or by no agent. (6) Target binary $t_{j}$: specifies whether $v_{j}$ coincides with agent $i$’s own target or with another agent’s target (if any). We further cluster all viewpoints with non-zero utility within a range $r_{b}$ into a beacon set [15], representing candidate regions to navigate. We further construct a planning graph $\bar{G}_{t}=(\bar{V}_{t},\bar{E}_{t})$ on $\mathcal{M}$, containing samples from both free and unknown cells, where edges connect collision-free pairs of neighbors. These structures allow us to define: (7) Navigation guidepost $g_{n,j}$: the Dijkstra trajectory on $\bar{G}_{t}$ from $v_{j}$ to the beacon nearest the agent’s assigned target. (8) Cooperation guidepost $g_{c,j}$: if any target remains unverified, i.e., not in the explored area, we compute the trajectory from $v_{j}$ to the beacon nearest the closest unverified target; otherwise $g_{c,j}=g_{n,j}$ if no targets remain.
@@ III-D2 Option Decoder $\pi_{\vartheta}$
During multi-agent navigation, an agent may either navigate toward its designated target or assist others by collecting more information that benefits the team, each demanding distinct and sometimes conflicting reasoning. In such cases, adaptive switching between self-directed and cooperative behaviors becomes crucial. To balance individual efficiency and collective performance for improved overall task success, we introduce the option-critic structure [1], where an option denotes a high-level behavioral mode characterized by a termination head function and an intra-option policy, while the critic provides a value-based estimation guiding high-level option/low-level action optimization. Specifically, we augment the policy network with an option decoder, which determines whether to continue the current high-level behavior or terminate and switch to another.
@@
MAContext w/o dual-stage [16] 522.07 ($\pm 162.09$) 64.59 510.11 ($\pm 170.3$) 62.64 495.39 ($\pm 151.08$) 61.03 583.66 ($\pm 177.65$) 71.52
MAContext [16] 503.3 ($\pm 152.98$) 61.63 495.45 ($\pm 129.99$) 61 478.49 ($\pm 110.82$) 59.16 500.77 ($\pm 168.78$) 60.53
EECBS [14] 526.12 ($\pm 151.22$) 66.08 501.78 ($\pm 127.77$) 62.64 477.12 ($\pm 129.86$) 59.49 468.3 ($\pm 136.55$) 58.6
LNS2 [13] 520.99 ($\pm 141.15$) 65.09 499.24 ($\pm 119.46$) 62.52 474.32 ($\pm 127.82$) 59.49 466.24 ($\pm 141.68$) 58.44
@@ III-E1 Multi-Agent Centralized Critic
Following MAAC [7], the critic computes a query $\mathbf{q}_{i}=W_{q}\mathbf{h}_{i}$ and attends to keys and values $\mathbf{k}_{j}=W_{k}\mathbf{u}_{j}$, $\mathbf{v}_{j}=W_{v}\mathbf{u}_{j}$ via
@@ III-E2 Option-Regularized Critic Learning
Our critic network adopts the same graph-transformer backbone as the policy network, but differs in its use of a privileged graph encoder. During training, this encoder has access to privileged ground-truth information (i.e., the ground-truth map), enabling the critic to provide low-variance estimations on long-term returns to the decentralized policies. Unlike the actor’s dual encoders (one for the combined graph and one for the current graph), the critic employs a single unified encoder that captures the complete relational context of all agents and their robot beliefs. Besides the state representation input, the critic also takes the option feature selected at the previous step as input, which benefits both option and waypoint optimization during critic learning [7]. The value estimate is explicitly conditioned on the agent’s previous option $z_{t-1}$, representing the high-level mode active at the current step. Specifically, the option embedding is added to the current node feature:
@@
On the other hand, following [5], the policy loss is further improved by our option-regularized policy network:
@@
Finally, the overall loss is $\mathcal{L}_{\text{policy}}=\mathcal{L}_{\pi}+\lambda_{\vartheta}\,\mathcal{L}_{\vartheta},$ where $\lambda_{\vartheta}$ balances termination learning. Following [5], the critic is optimized via a temporal-difference (TD) regression objective.
@@ IV-A Comparisons in Simulated Maps
Inspired by recent navigation and exploration benchmarks [16, 23], we introduce a new benchmark comprising hundreds of large, maze-like, and challenging simulated maps, in which we evaluate ORION against representative baselines under partially known environments. For the learning-based baseline, we consider MAContext [16], a multi-agent extension of a DRL navigation planner designed for partially known environments. For classical MAPF solvers, we include 1) LNS2 [13], a highly efficient and scalable planner that iteratively repairs collision-containing path sets by replanning for a subset of colliding agents, and 2) EECBS [14], a bounded-suboptimal MAPF planner that combines Explicit Estimation Search at the high level with focal search at the low level. Since both LNS2 and EECBS operate on static maps, we deploy them in our partially known setting by triggering replanning whenever new changes are observed by any agent in the environment. To further analyze the contribution of our design, we additionally evaluate two ablation variants: 1) ORION w/o option, removing our option-critic structure, and 2) ORION w/o dual-stage, disabling our pre-/post-arrival stages. These variants help isolate the role of our proposed option-critic’s hierarchical decision-making and dual-stage cooperation strategy.